Grandma, Here’s How to Use Your New Computer

William Stein April 2001

Grandma, congratulations on your new laptop!

Contents 1 Setting the Laptop Up 2 2 Turning the Laptop Off and On 3

2.1 Turning it On . . . . . . . . . . . . . . . . . . 3 2.2 Turning it Off . . . . . . . . . . . . . . . . . . 5

3 Using the Track Pad to Move the Pointer 6 4 Connecting to and Disconnecting from the

Internet 7 4.1 Connecting . . . . . . . . . . . . . . . . . . . . 8 4.2 Disconnecting . . . . . . . . . . . . . . . . . . 9 4.3 Your Internet Service Provider . . . . . . . . . 9

5 Using Email 10 6 Viewing Web Pages 12 7 What to do if the computer “Freezes”, or

doesn’t resume correctly. 13 8 Playing Games 13 9 Playing Music Built Into the Computer 14 10 Playing CDs 16 11 Technical Information 16 1 Setting the Laptop Up

1. This picture illustrates how to plug everything in:

2. The laptop should usually be plugged into the wall, because the built-in battery doesn’t last very long (15 minutes, maybe). To plug the laptop into the wall, plug one connector of the power cord into the right back side of the laptop, and the other into the wall. Don’t worry if you accidently unplug the laptop for a few minutes while

it is on.

2 Turning the Laptop Off and On

2.1 Turning it On

• If the laptop is off, do the following.

1. Slide the on switch to the right (see the illustration below).

2. Wait about 1 minute for the login: prompt.

3. Type grandma at the prompt. (If you mess up, it will ask you for a password. Just press Enter, and wait for the login: prompt to re-appear.

4. Wait 1 minute 20 seconds for the graphical window system to appear.

5. Now you are ready to work.

6. Notice the button accross the bottom of the screen:

(a) The one that says “K” brings up a large menu.

(b) The one that looks like a plus activates your modem.

(c) The big “E” runs your EMAIL program.

(d) The globe and wheel are used to bring up progams

that let you surf the Internet

(e) The cards icon starts a video Poker game.

(f) The notes start a Music player.

• If the laptop is suspended, do the following:

1. Open the screen.

2. Press the round “Moon Button” on the left near the screen (see illustration below).

3. The screen should appear, and you are ready to work.

4. If this doesn’t work, maybe you accidently unplugged the laptop for a long time and the battery died, or the power went out for a while. Follow the off steps above, except step 2 will take about 5 minutes instead of 1 minute.

2.2 Turning it Off

• To suspend the computer:

1. Simply close the lid. That’s all.

• To turn the computer off, so that you can unplug it:

1. Click the K in the lower left corner of the screen to bring up the menu.

2. Click on Logout, which is the menu option right above the K.

3. Click Logout again in the pop-up dialog.

4. Wait at least 1 minute for everything to properly shut down.

3 Using the Track Pad to Move the Pointer

These days computer interfaces are typically very graphical. In order to work, there is a little pointer that you have to move around on the screen. Unfortunately, you can’t just touch the screen (screens with touch sensitivity usually don’t look so good). Instead, you move your finger around on the “track pad”, which is that black pad with two little buttons in the front of the laptop.

You put your index finger on the pad and slide it. You’ll have to practice. Dennis will show you how.

4 Connecting to and Disconnecting from the Internet

The Internet is an interconnected network of probably hundreds of millions of computers. When you are connected to the Internet, When the laptop is connected to the Internet, you can almost instantly send messages to mom, Dennis, and I, and millions of other people, you can read the news, and do many other useful things. In order to use the Internet, your computer has to be plugged into the US Robotics modem that I sent you, and the modem has to be plugged into your phone line. Also, the modem AC adapter has to plugged in to a power outlet.

4.1 Connecting

To connect to the internet, make sure you do the following:

1. Make sure the modem is ON (the switch is towards the I, not the O) and PLUGGED IN (at least one red light is on).

2. Click on the icon that is second from the lower left and looks

3. Your modem will make loud and crazy sounding buzzing sounds for about 30 seconds.

4. Congratulations, if several of the modem lights are on, you are now connected to the Internet!

5. It’s conceivable that, in fact, something went wrong and you aren’t connected. If so, click the connect button again and wait.

6. If you still can’t connect, give up and try later.

7. If you still can’t connect, get in touch with me.

4.2 Disconnecting

You have many choices:

• Turn the modem off and on.

• Shut the lid of the computer to suspend the computer (see Section 2.2).

• Turn the computer off (see Section 2).

4.3 Your Internet Service Provider

The web page of your ISP (Internet Service Provider) is

http://www.aplus.net

Your account manager is

Abel Fazel Poor
ABACUS America Inc. d.b.a A+Net
5266 Eastgate Mall
San Diego, CA 92121
(868) 558-8449 x 101 or
1-877-275-8763 x 101

The technical support number is 858-410-6900. If something goes wrong and you can’t connect to the web, it’s probably best to ask me to contact them, as it’ll be impossible for you to explain to them how your computer connects to the web.

You have a 20 month pre-paid ($299) account. So you shouldn’t worry about paying them anything until at least Christmas of 2002.

Your account information is:

username: amp/grandma
password: roxy3584                  // secret!!

email login: grandma

Access numbers

Cambridge: (617)8125599
Tucson: (520)6292980
Flagstaff: (520)213-3320

5 Using Email

Click on the button on the bottom of the screen that looks like a big “E” with an envelope next to it. It is the third button from the left. Wait at least 10 seconds.

The email program is divided into three windows. The upper-left window lists your folders; click on a folder to select it. The upper-right window lists the messages in the folder selected to the left; click on a message to view it. Messages are viewed in the window that goes across the bottom of the screen.

• To Get New Mail:

1. Make sure you are connected to the web, i.e. that lots of lights on your modem are ON. If you’re not connected to the web, do so now (see Section 4).

2. Click on the button that looks like a tray at the top of the email program. It is the fourth button from the left. A status bar in the lower right will appear. If somebody like Mom sends you a big message, it may take a while to download. If it is taking forever to download, click the “X” next to the status bar and then reconnect to the internet as described in Section 4.

3. Read your mail!

• To Respond to a Message:

1. While viewing the message you want to respond to, push the “r” key on the keyboard.

2. Use the down arrow on the lower-right part of the keyboard to move the blinking “cursor” to the main text of the message. (I.e., Hit the down arrow three times.)

3. Start typing. You’ll see the text of the message you’re responding to quoted in the same editor window. This is also sent, so the recipient will know what you’re responding to.

4. When you’re done, click the send button. This is the button in the upper left corner of the window that you’re typing the message in. Alternatively, if that button disappears, select Send from the Message menu. Alternatively, if you don’t know how to do that, hold the little key labeled Ctrl down (lower left corner of keyboard) and press the Enter key.

5. Hopefully the message is sent off, over the phone line.

6. If you get an error message, it almost certainly means that for some reason your modem is connected to the internet. See Section 4. Reconnect your modem, then select Send Queued from the File menu.

There are many other things that you can do with the email program. However, getting new mail and responding to mail is all you need to start talking with us.

6 Viewing Web Pages

Make sure you are connected to the Internet (see Section 4).

1. You have two web browsers. To start the first (which I recommend), click on the fourth button from the left, which looks sort of like Earth. To start the second (which I don’t recommend), click on the fifth button from the left at the bottom of the screen, which looks like a wheel.

2. You can type an address into the blank at the very top of the web browser and hit return to “go there”. For example,

if you type modular.fas.harvard.edu and hit return,

you’ll be directed to my web page. You can click on links on each web page to see new web pages.

3. You can also type words into the search box at the middle of the startup screen to “search the web”.

7 What to do if the computer “Freezes”, or doesn’t resume correctly.

If the computer completely stops responding to your commands for a few minutes, DON’T PANIC. Do the following.

1. Turn the computer off, as follows:

• Hold the button with the moon above it at the front down

• Move the lever to the right, with the moon button held down.

• Wait a few seconds.

2. Turn the computer back on by sliding the lever to the right.

3. Wait. The computer was not shut down properly, so it will have to check the disk for any errors. This will take from 5 to 10 minutes. Be patient.

4. When the login: prompt appears, type grandma and press the Enter button. (If the login prompt doesn’t appear, get in touch with me...)

5. Wait about 3 minutes as the window system loads.

6. Everything is fixed.

8 Playing Games

Your laptop comes with lots of card and other games. You don’t have to be connected to the Internet to use these games. For example, click on the “Poker” icon, which is 6th from the left at the bottom of the screen in order to play (free) video poker. The are also a few dozen other games available as follows.

1. Click on the “K” in the lower left corner of the screen.

2. A menu will appear.

3. Click on “Games” in the menu.

4. Select a game, such as “patience’.

5. WAIT for it to load.

6. Have fun.

9 Playing Music Built Into the Computer

Click on the icon that looks like notes in the middle bottom of the screen to play the music builtin to the laptop. There are 326 songs available, over 21 hours of good music, which you can see on the right hand side, and click to listen to any particular song. The volume control is a dial on the left side of the computer.

The builtin speakers aren’t very loud, but you could get external speakers. WARNING: Playing music can slow the computer down a lot.

10 Playing CDs

First there were records, then 8-tracks, then cassettes, then CDs, which are round shiny discs that contain about an hour of music. Your laptop has a builtin CD player.

To play a CD:

1. Insert the CD.

 2. Click [K] ⟶ Multimedia ⟶ CD Player.

11 Technical Information

Your laptop has 48MB of RAM, a 10GB hard drive (huge!), a 133 Mhz Pentium MMX processor, and an active matrix 800x600 display. The modem is a US Robotics 56K external modem. The laptop is running the RedHat Linux 7.0 operating system along with the KDE 2.1 desktop environment.